Right triangle ABC has one leg of length 7cm and one leg length 11 cm. determine the length of the hypotenuse, the perimeter of the triangle and the area of the triangle.
Answer by mananth(16949)

one leg 7cm
other leg 11 cm.
..
hypotenuse^2= 7^2+11^2
H^2=49+121
H^2=170
H= sqrt 170
Hypotenuse = 13.03 cm
..
Perimeter = sum of all sides
=7+11+13.03=31.03 cm
..
Area of right triangle = 1/2 *one leg * other leg
=1/2 * 7*11
38.5 cm^2
..